OpenStudy (anonymous):

What number is 30% of 140?

OpenStudy (ja1):

140 x 0.7

OpenStudy (anonymous):

30% is equal to 30/100 or 3/10.

OpenStudy (mathstudent55):

% means 1/100 or 0.01 "of" in math ususally means times 30% of 140 means 30% * 140 = 30 * 1/100 * 140 = 0.3 * 140

OpenStudy (ja1):

wait my bad i meant .30 lol not .7

OpenStudy (ja1):

So it would be 140*.30

OpenStudy (anonymous):

42??

OpenStudy (ja1):

Yes

OpenStudy (ja1):

140*.30 = 42

OpenStudy (anonymous):

Okay, I hope that right :) Thank you (:

OpenStudy (ja1):

Always a pleasure to help :)

OpenStudy (anonymous):

Great